OpenStudy (eva12):

1) find all positive integers solutions of phi(n)=12

OpenStudy (eva12):

I know the divisors of 12 :1,2,3,4,6,12. Possible factors :2,3,4,5,7,13 how do I find possible solutions?

OpenStudy (vats-sagar):

phi(n) ?

OpenStudy (anonymous):

its the totient function right?

OpenStudy (eva12):

Yes

OpenStudy (anonymous):

one solution will be 13

OpenStudy (eva12):

Is that since 13-1 =12 => 12*1 since phi (13)* phi(2) =12*1 right?

OpenStudy (anonymous):

yes for a prime number say p\[\phi(p)=p-1\]and 13 satisfies the case but there can be more solutions

OpenStudy (anonymous):

when n was prime the solution we got was 13 :) prime factorization of n if (n is not prime of course)->\[n=p_1^{x_1} \times p_2^{x_2} \times p_3 ^{x_3}......\]where \(p_1,~p_2,...\) are primes and \(x_1,~x_2,.... \ge 1\) now we know that totient function is multiplicative so ->\[\phi(n)=\phi(p_1^{x_1}) \times \phi(p_2^{x_2}).....\]since \(p_1,~p_2,... ~are~primes~we~can~write~it~as->\)\[\phi (n)=\left( p_1^{x_1}-p_1^{x_1-1} \right)\left(p_2^{x_2}-p_2^{x_2-1} \right)......\] \[12=\left( p_1^{x_1}-p_1^{x_1-1} \right)\left(p_2^{x_2}-p_2^{x_2-1} \right)......\] we know that 12=1x12 =2x6 =3x4 so we can have only 3 cases when n is not prime now you just have to evaluate these cases and then u can get more solutions from here

OpenStudy (loser66):

It is a very long solution for this. ha!! But let me try \(n =\prod_{i=1}^s p_i^{a_i}\), \(\phi (n) = \prod_{i=1}^s p_i^{a_i -1} (p_i-1)\) Hence \((p_i -1)| \phi (n)=12\) . Therefore, if \(p_i > 17, p_i-1 =16 > 12\rightarrow 16\cancel | 12\) Hence \(p_i \leq 13\) then \(\large n = 2^a~ 3^b~ 5^c~ 7^d ~11^e ~13^f\) Moreover, \(p_i^{a_i} | n\) , hence 1) \(2^{a-1} |12 \implies (a-1)\leq 2, hence~~ a =0,1,2,3 \) 2) \(3^{b-1}| 12\implies b-1\leq1,~hence~ b=0,1,2\) 3) \(5^{c-1}|12\implies c-1 \leq 0~hence c =0,1 \) 4) \(7^{d-1}|12\implies d-1\leq 0~hence~d=0,1\) 5) \(11^{e-1}|12\implies e-1\leq 0~hence~e =0,1\) 6)\(13^{f-1}|12\implies f-1\leq 0, ~~hence~f =0,1\) ha!!! now combine all If a =b=c=d=e= 0 \(f\neq 0\) , that is f =1, hence n = 13, \(\phi(13) =12\color{red}{\checkmark}\) If a=b=c =d =f =0, \(e\neq 0\) , that is e =1, hence n = 11, but \(\phi (11)\neq 12\) reject!! If a=b=c=e=f =0 \(d \neq 0\)........ You do all of the options, the final result are n = 13,21,26,28, 36,42 Good luck.

OpenStudy (eva12):

Thank you